OpenStudy (anonymous):

Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x2 - 4x y = -2x2 + 8x - 12

OpenStudy (anonymous):

Its a test problem im having trouble on :C

OpenStudy (mathmate):

Equation 1: 1. concave up. 2. vertex at (2,-4) 3. y-intercept = 0 4. zeroes at x=0 and x=4. I'll leave it to you to do the second equation for the compare and contrast.

OpenStudy (anonymous):

how did you get that tho

OpenStudy (mathmate):

Equation 1: y = x2 - 4x 1. concave up. [leading coefficient is positive] 2. vertex at (2,-4) [rewrite in canonical form as y=(x-h)^2+k =(x-2)^2-4, then (h,k) is the vertex.] 3. y-intercept = 0 [from the general form ax^2+bx+c, where c=0 => y-intercept=0] 4. zeroes at x=0 and x=4. [rewrite y=x^2-4x=x(x-4) => zeroes at x=0 and x=4].